# Single-shot thermometry of simulated Bose–Einstein condensates using artificial intelligence

### v1

@@ 6 Summary
We have introduced a proof-of-principle machine learning model which can, within a single shot, estimate the values of the chemical potential and temperature of a Bose-condensed cloud of atoms. We use convolutional neural networks — the foundation of most image recognition models — to take an atomic density profile and estimate important thermodynamic parameters. We have demonstrated that our model for a harmonically trapped condensate can estimate to some degree the chemical potential and temperature for systems that it has not previously been trained on, including during thermalisation, and on toroidally trapped condensates. This work joins recent applications of machine learning in the quantum fluids literature, such as the identification of topological defects such as vortices and solitons, and the reconstruction of vortex filaments [19, 20, 21].
@@ 8 Acknowledgements
We thank the EPSRC, Grant No. EP/T015241/1 and EP/T518001/1, for funding this project.
@@ Appendix F Visualisation of the second and third convolutional layers
Figure 11: The cross-correlation, $\zeta^{1}$, of a single-shot atomic density, $\rho(x,y)$, with 12 weights kernels, $\mathsf{W}^{1}_{i}$, $i\in\{1,\cdots,12\}$, as determined by equation (15). The cross-correlation may result in positive or negative values, as indicated by the colour bars.

### v2

@@ 6 Summary
We have introduced a proof-of-principle machine learning model which can, with a single image, estimate the values of the chemical potential and temperature of a Bose-condensed cloud of atoms. We use convolutional neural networks — the foundation of most image recognition models — to take an atomic density profile and estimate important thermodynamic parameters. We have demonstrated that our model for a harmonically trapped condensate can estimate to some degree the chemical potential and temperature for systems that it has not previously been trained on, including during thermalisation, and on toroidally trapped condensates. This work joins recent applications of machine learning in the quantum fluids literature, such as the identification of topological defects such as vortices and solitons, and the reconstruction of vortex filaments [19, 20, 21].
@@ 8 Acknowledgements
We thank K. E. Wilson for insightful discussions about the experimental considerations of this work. We thank the EPSRC, Grant No. EP/T015241/1 and EP/T518001/1, for funding this project.
@@ Appendix F Visualisation of the second and third convolutional layers
Figure 11: The cross-correlation, $\zeta^{1}$, of a single image of an atomic density, $\rho(x,y)$, with 12 weights kernels, $\mathsf{W}^{1}_{i}$, $i\in\{1,\cdots,12\}$, as determined by equation (15). The cross-correlation may result in positive or negative values, as indicated by the colour bars.